# Uncertainty Quantification for LLMs through Minimum Bayes Risk: Bridging Confidence and Consistency

### v4

@@ 5.1 Experimental Setup
To evaluate the effectiveness of the proposed method, we extended the LM-Polygraph library (Vashurin et al., 2025; Fadeeva et al., 2023) by implementing our approach within its framework. Since the library already includes tools for calculating other uncertainty scores, it provided a convenient and efficient environment for setting up and running experiments. The primary objective of our experiments is to evaluate whether our method offers improved performance in key tasks such as question answering (QA), summarization (SUM), and machine translation (MT), compared to existing baselines.
@@
Datasets. For QA, we selected diverse datasets to capture a variety of challenges: TriviaQA (Joshi et al., 2017), an open-domain factual QA dataset; CoQA (Reddy et al., 2019), a conversational QA benchmark requiring multi-turn contextual understanding; MMLU (Hendrycks et al., 2021), a multi-task dataset spanning 57 topics to test broad knowledge; and GSM8k (Cobbe et al., 2021), which focuses on grade-school math problems requiring logical reasoning. For translation, we evaluated our method on WMT14 French-English (Bojar et al., 2014) and WMT19 German-English (Barrault et al., 2019). Finally, for summarization, we used XSUM (Narayan et al., 2018), a dataset of complex documents paired with concise abstractive summaries.
For all datasets, we follow (Vashurin et al., 2025) for selecting the subsets, for prompt formatting, and for the number and sourcing of few-shot examples.
@@
Models. We evaluated our method using the base versions of three open-weight language models: LLaMA 3.1 8B Touvron et al. (2023), Mistral 7B Jiang et al. (2023), and Falcon 3 7B Team (2024). The open-source nature of these models provides full access to their token probabilities, which are essential for implementing our UQ method. For all models, we consider base versions, without instruction tuning.
@@
Similarity Function. To measure the similarity, we use the RoBERTa-large cross-encoder model, fine-tuned on the Semantic Textual Similarity benchmark dataset (Liu et al., 2019; Reimers and Gurevych, 2019; Cer et al., 2017). This model is widely regarded as one of the most reliable and commonly used approaches for evaluating sentence similarity. The cross-encoder processes two sequences jointly and directly outputs a similarity score ranging from 0 to 1, providing a nuanced measure.
Appendix C.1 contains comparative experiments with cross-encoder and other choices of the similarity function, substantiating this choice.
@@
Baselines. We compare the performance of the proposed method against a diverse set of baselines and state-of-the-art UQ scores, including confidence-based, consistency-based, and hybrid approaches. Information-based methods include Maximum Sequence Probability (MSP), Perplexity (PPL), Mean Token Entropy (MTE), Monte Carlo Sequence Entropy (MCSE), and Monte Carlo Normalized Sequence Entropy (MCNSE). Consistency-based methods include the Degree Matrix (DegMat) and the Sum of Eigenvalues of the Graph Laplacian (EigValLaplacian). Finally, hybrid methods include Semantic Entropy and SAR. All formulations for these baselines can be found in Appendix D. Lastly, we evaluate performance of (Consistency) and (Consitency Light) as an uncertainty measure.
@@
Evaluation measure. As our evaluation measure, we choose the Prediction Rejection Ratio (PRR), which measures the effectiveness of the uncertainty scores for identifying high-quality predictions (Malinin and Gales, 2021). PRR operates by progressively rejecting predictions with uncertainty scores above a threshold and observing how the average quality of the remaining predictions changes (Figure 3). The metric is calculated as the ratio of two areas: the area between the Prediction Rejection (PR) curves for the evaluated uncertainty score and a random baseline, and the area between the oracle (the ideal uncertainty score that perfectly ranks instances by quality) and the random baseline. Formally, PRR is defined as follows:
@@
Higher PRR values indicate better alignment of uncertainty scores with prediction quality, approaching the performance of an oracle. To ensure practical applicability, we compute PRR only up to a rejection threshold of 50%, preventing cases where excessive rejection artificially inflates quality measures.
For some of the datasets we also report AUROC results, see Appendix G.
@@
Quality Measures. PRR requires an appropriate quality measure for each specific task to effectively evaluate the model output. For question-answering tasks, we use Accuracy to directly evaluate whether the generated answers match the ground truth in short-form QA tasks (e.g., MMLU), and we use the AlignScore between the correct answer and generated sequence for assessing the performance for long-form QA tasks (Zha et al., 2023). For summarization tasks, we use AlignScore to measure the alignment between the output summary and the input document. It serves as a quality indicator by evaluating the relevance and the overlap between the generated content and the source text. For translation tasks, we use COMET, as it captures both semantic adequacy and fluency, ensuring that translations are accurate and linguistically appropriate (Rei et al., 2020).
@@
To further improve comprehensiveness of reported results, we considered alternative choices of quality measures, and report corresponding PRR scores in Appendix F.
@@
Generation Setup. We discuss the generation parameters, the decoding strategy, and sample selection procedure in depth in Appendix A. In short, we report evaluation results in two distinct setups: greedy decoding and stochastic sampling with a focus on the most probable sequence among the generated outputs (best-sample). These two setups offer the highest-quality outputs and are the most reasonable generation approaches in practice.
@@
CoCoA Light Details. The multilayer perceptron is used as an auxiliary model. The features used are embeddings from the middle layers of the base LLM, which are the most informative according to recent work Chen et al. (2024). In particular, for Llama 3.1 8B and Mistral 7B models, embeddings are taken from the 16th layer, and for Falcon 3 7B from the 14th layer.
@@ 5.2 Results
Main results. Table 1 shows the PRR scores under the greedy generation setup (See Appendix B for Best Sample and MBR Sample). We report aggregated PRR for each type of task – question answering, neural machine translation (NMT), and summarization (SUM) – by averaging the results across all relevant datasets (e.g., TriviaQA, MMLU, CoQA, GSM8k for QA). This aggregated score provides a concise measure of the performance for each model for each task. Detailed results for each dataset separately can be found in Appendix E.
@@
Ablation study. As a part of our ablation study, we investigate the impact of different similarity measures (see Appendix C.1). We find that for some tasks, the similarity score computed by the Cross-encoder does not yield optimal performance. For example, for question-answering tasks on CoQA and TriviaQA, NLI-derived similarity performs better than the Cross-encoder similarity and outperforms the pure consistency-based uncertainty discussed above.
@@
Lastly, in Appendix C.2, we also consider alternative formulations of the information-based metric that do not rely on logarithmic transformations. While we primarily use logarithms due to their numerical stability, we explore an alternative approach by converting these values back to probabilities and analyzing their impact on uncertainty estimation. Our findings indicate that both formulations exhibit consistent performance and yield similar results. This suggests that while logarithmic transformations enhance numerical stability, the choice between log-based and probability-based formulations does not affect much the overall performance.
@@ 6 Limitations
Limited Sample Size. CoCoA estimates consistency by sampling multiple outputs. Generating many samples is computationally expensive, so our experiments—like most sampling-based methods—use relatively small sets. While even a few samples can yield meaningful estimates, they may miss the full diversity of model outputs for complex prompts.
@@ 7 Conclusion
Moving forward, several directions are open for further exploration. These include incorporating more adaptive sampling strategies that efficiently capture the model output space, refining semantic similarity functions for domain-specific tasks, and improving calibration techniques to strengthen the confidence metrics of the model.

### v5

@@ 5 Experiments
In this section, we present the experimental setup, the results, and the ablations.
@@ 5.1 Experimental Setup
To evaluate the effectiveness of our proposed framework, we extended the LM-Polygraph library (vashurin2024benchmarkinguncertaintyquantificationmethods; fadeeva-etal-2023-lm). Since it already includes tools for calculating other uncertainty scores, it provided a convenient and efficient environment for setting up and running experiments. The primary objective of our experiments is to evaluate whether our method offers improved performance in key tasks such as question answering (QA), text summarization (SUM), and machine translation (MT), compared to existing baselines.
@@
Datasets. For QA, we selected diverse datasets to capture a variety of challenges: TriviaQA (joshi-etal-2017-triviaqa), an open-domain factual QA dataset; CoQA (coqa), a conversational QA benchmark requiring multi-turn contextual understanding; MMLU (mmlu), a multi-task dataset spanning 57 topics to test broad knowledge; and GSM8k (gsm8k), which focuses on grade-school math problems requiring logical reasoning. For translation, we evaluated our method on WMT14 French-English (wmt14) and WMT19 German-English (wmt19translate). Finally, for summarization, we used XSUM (xsum), a dataset of complex documents paired with concise abstractive summaries. For all datasets, we follow (vashurin2024benchmarkinguncertaintyquantificationmethods) in subset selection, prompt formatting, and few-shot example sourcing.
Models.
@@
We evaluated our method on the base versions of three open-weight language models: LLaMA 3.1 8B (touvron2023llama), Mistral 7B (mistral), and Falcon 3 7B (Falcon3). The open-weight nature of these models enables direct access to token probabilities, which is crucial for implementing our UQ method. All experiments were conducted using the base (non–instruction-tuned) variants. We additionally report results on the larger Gemma 3 12B-Base model (gemmateam2025gemma3technicalreport), with detailed analysis provided in Appendix J.
Similarity Function.
@@
To measure the similarity between two generations, we use the RoBERTa-large cross-encoder model, fine-tuned on the Semantic Textual Similarity benchmark dataset (liu2019robertarobustlyoptimizedbert; DBLP:journals/corr/abs-1908-10084; huggingface:dataset:stsb_multi_mt). This model is widely regarded as one of the most reliable and commonly used approaches for evaluating sentence similarity. The cross-encoder processes two sequences jointly and directly outputs a similarity score ranging from 0 to 1, providing a nuanced measure. Appendix C.1 contains comparative experiments with cross-encoder and other choices of the similarity function, substantiating this choice.
Baselines.
@@
We compare the performance of the proposed method against a diverse set of baselines and state-of-the-art UQ scores, including confidence-based, consistency-based, and hybrid approaches. For information-based approaches, we evaluate Sequence Probability (SP), Perplexity (PPL), Mean Token Entropy (MTE), Monte Carlo Sequence Entropy (MCSE), and Monte Carlo Normalized Sequence Entropy (MCNSE). In the consistency-based category, we consider the Degree Matrix (DegMat) and the Sum of Eigenvalues of the Graph Laplacian (EigValLaplacian). Finally, we include hybrid methods Semantic Entropy and SAR, as well as verbalized confidence method P(true) (kadavath2022language). All formulations for these baselines can be found in Appendix D. Finally, we evaluate the performance of (Consistency) and (Consistency Light) as an uncertainty measure.
Evaluation Measure.
@@
As our evaluation measure, we chose the Prediction Rejection Ratio (PRR), which measures the effectiveness of the uncertainty scores for identifying high-quality predictions (malinin2020uncertainty). PRR operates by progressively rejecting predictions with uncertainty scores above a threshold and observing how the average quality of the remaining predictions changes (see Figure 3). It is calculated as the ratio of two areas: the area between the Prediction Rejection (PR) curves for the evaluated uncertainty score and a random baseline, and the area between the oracle (the ideal uncertainty score that perfectly ranks instances by quality) and the random baseline. Formally, PRR is defined as follows:
@@
Higher PRR values indicate better alignment of uncertainty scores with prediction quality, approaching the performance of an oracle. To ensure practical applicability, we compute PRR only up to a rejection threshold of 50%, preventing cases where excessive rejection artificially inflates the quality measures.
For the QA datasets, we further report AUROC (see Appendix G).
Quality Measures.
@@
PRR requires an appropriate quality measure for each specific task to effectively evaluate the model output. For question–answering tasks, we use Accuracy to directly evaluate whether the generated answers match the ground truth in short-form QA tasks (e.g., MMLU), and we use the AlignScore between the correct answer and generated sequence for assessing the performance for long-form QA tasks (zha2023alignscore). For summarization tasks, we use AlignScore to measure the alignment between the output summary and the input document. It serves as a quality indicator by evaluating the relevance and the overlap between the generated content and the source text. For translation tasks, we use COMET, as it captures both semantic adequacy and fluency, ensuring that translations are accurate and linguistically appropriate (rei-etal-2020-comet).
@@
To further improve the comprehensiveness of the reported results, we considered alternative choices of quality measures, and report the corresponding PRR scores in Appendix F.
Generation Setup.
@@
We discuss the generation parameters, the decoding strategy, and the sample selection procedure in depth in Appendix A. In short, we report the evaluation results in two distinct setups: greedy decoding and stochastic sampling with focus on the most probable sequence among the generated outputs (best-sample). These two setups offer the highest-quality outputs and are the most reasonable generation approaches in practice.
CoCoA Light Details.
@@
We use a multilayer perceptron network as an auxiliary model. The features used are embeddings from the middle layers of the base LLM, which are the most informative according to recent work (chen2024inside). In particular, for Llama 3.1 8B and Mistral 7B models, we take embeddings from the 16th layer, and for Falcon 3 7B from the 14th layer. More details on the training are available in Appendix I.
@@ 5.2 Results
Table 1 shows the PRR scores under the greedy generation setup (See Appendix B for Best Sample and MBR Sample). We report aggregated PRR for each type of task – question answering, neural machine translation (NMT), and summarization (SUM) – by averaging the results across all relevant datasets (e.g., TriviaQA, MMLU, CoQA, GSM8k for QA). This aggregated score provides a concise measure of the performance for each model for each task. Detailed results for each dataset separately can be found in Appendix E.
@@
5.3 Ablations
Similarity Function.
We investigate the impact of different similarity measures (see Appendix C.1). On average, the cross-encoder provides strong performance. However, for summarization tasks, AlignScore yields better results, while in long-form QA, similarity based on Natural Language Inference (NLI) sometimes outperforms the cross-encoder. We hypothesize that for generation with shorter outputs (1–2 sentences), any capable NLI or cross-encoder model is sufficient. For longer generations, approaches that estimate similarity over chunks and then aggregate scores (e.g., AlignScore) may be more effective. Nevertheless, across all tasks, the cross-encoder achieves the best overall performance and, in scenarios where tuning is not feasible, serves as a strong default choice.
Alternative Formulations.
@@ 5.3 Ablations
Lastly, in Appendix C.2, we also consider alternative formulations of the information-based metric that do not rely on logarithmic transformations. While we primarily use logarithms due to their numerical stability, we explore an alternative approach by converting these values back to probabilities and analyzing their impact on uncertainty quantification. Our findings indicate that both formulations exhibit consistent performance and yield similar results. This suggests that while logarithmic transformations enhance numerical stability, the choice between log-based and probability-based formulations does not affect much the overall performance.
@@ 6 Limitations
Limited Sample Size. CoCoA estimates consistency by sampling multiple outputs. Generating many samples is computationally expensive, and thus our experiments, e.g., most sampling-based methods, use relatively small sets. While even a few samples can yield meaningful estimates, they may miss the full diversity of model outputs for complex prompts.
@@ 7 Conclusion and Future Work
Moving forward, several directions are open for further exploration. These include incorporating more adaptive sampling strategies that efficiently capture the model output space, refining the semantic similarity functions for domain-specific tasks, like code generation or commonsense reasoning.
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
Answer: [Yes]
Justification: The abstract and introduction accurately describe the proposed family of Confidence and Consistency-based Approaches (CoCoA) to UQ. Theoretical contributions include new methods that merge information- and consistency-based measures for UQ in LLMs. Experimental validation of the new approaches is done across different NLP tasks: question answering, summarization, and translation. It aligns with the content presented in the paper (Section 3, 5.2, Appendix B).
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: Section 6 discusses limitations, including relying on an information-based confidence score and a semantic similarity function, having limited sample sets, and depending on a quality metric.
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory assumptions and proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [N/A]
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental result reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Justification: Section 5.1 outlines the experimental setup, datasets and evaluation metrics. The description of datasets used for each task includes TriviaQA, CoQA, MMLU, GSM8k for Question Answering, WMT14 French-English, WMT19 German-English for machine translation, and XSUM for summarization. The models used: LLaMA 3.1 8B, Mistral 7B and Falcon 3 7B. The similarity function is specified as RoBERTa-large fine-tuned on the Semantic Textual Similarity benchmark. The evaluation metric, Prediction Rejection Ratio (PRR), is clearly defined, and task-specific quality measures such as Accuracy, AlignScore, and COMET are used. Furthermore, the training setup for the CoCoA Light method, including the use of multilayer perceptron and middle layer embeddings of the base LLM, is described. The appendix provides additional details on ablation studies, similarity metrics, and decoding strategies (Appendix C and E).
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [Yes]
Justification: Code is open source and is provided as an extension to the LM-Polygraph library (Section 5.1).
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental setting/details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Justification: Experimental settings, including model architecture (base LLaMA 3.1 8B, Mistral 7B, Falcon 3 7B), embedding layer selection for auxiliary training (16th or 14th layer depending on model), and the auxiliary model structure (multilayer perceptron) are described in Section 5.1 and Appendix A.
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment statistical significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [No]
Justification: The paper does not provide errorbars as it operates in (almost) deterministic scenario: we operate with pre-trained models and there is no training involved whatsoever. The only randomization involved is the one related so sampling multiple sequences for consistency-based methods. The choice to not compute errorbars here is deliberate as our inference involves heavy computations and it is not possible to conduct main experiments multiple times due to that. However, we conducted experiments on 3 different LLMs and 7 datasets, which is equivalent to 21 independent run, demonstrating the robustness and reliability of our findings. For this reason, we do not report error bars, as the results are stable.
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments compute resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: Appendix H discusses computational resources used to produce experimental results.
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code of ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: Our research does not involve human subjects or ethically sensitive applications. We believe it conforms the NeurIPS Code of Ethics.
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [Yes]
Justification: Section 1 discusses the importance of Uncertainty Quantification, highlighting its positive societal impacts (helps to reduce overreliance on automated systems and encourages a more responsible LLM usage). We do not foresee any negative negative societal impact of this work.
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [N/A]
Justification: N/A.
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: The paper properly cites the original sources of models (e.g., LLaMA 3.1 8B, Mistral 7B, Falcon 3 7B), datasets (CoQA, TriviaQA, MMLU, GSM8k, XSUM, WMT14FrEn, WMT19DeEn), and tools (e.g., RoBERTa, LM-Polygraph). However, it does not explicitly mention the specific licences and terms of use for these assets.
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [No]
Justification: The paper introduces a new family of uncertainty quantification methods called CoCoA (Confidence and Consistency-based Approaches) and an approximated variant, CoCoA Light. However, it does not release any new dataset, model, or a software tool alongside the paper, and does not provide accompanying documentation or artifacts.
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and research with human subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [N/A]
Justification: The research does not involve crowdsourcing experiments or research with human subjects.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional review board (IRB) approvals or equivalent for research with human subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [N/A]
Justification: The research does not involve crowdsourcing experiments or research with human subjects.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
16.
Declaration of LLM usage
Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
Answer: [Yes]
Justification: The paper centers its contributions on the use of Large Language Models (LLMs) as the core subject of study. It introduces new methods for uncertainty quantification in LLMs, specifically evaluating base models such as LLaMA 3.1 8B, Mistral 7B, and Falcon 3 7B. Apart from that LLMs were used solely to aid in editing the final manuscript and writing some of the visualization code.
Guidelines:
The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.